OpenStudy (nudeoua):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) greater than or equal to −12? x greater than or equal to −5 x less than or equal to −5 −5 less than or equal to x A number line with a closed circle on negative 5 and shading to the right

imqwerty (imqwerty):

\[2(x-1) \ge -12\]\[(x-1) \ge -6\]\[x \ge -5\] clearly x is greater than or equal to -5 the option is is wrong cause it says the opposite

imqwerty (imqwerty):

*option B

OpenStudy (nudeoua):

Thank you!